# HIP-LLM: A Hierarchical Imprecise Probability Approach to Reliability Assessment of Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly deployed across diverse domains, raising the need for rigorous reliability assessment methods. Existing benchmark-based evaluations primarily offer descriptive statistics of model accuracy over datasets, providing limited insight into the probabilistic behavior of LLMs under real operational conditions. This paper introduces HIP-LLM, a Hierarchical Imprecise Probability framework for modeling and inferring LLM reliability. Building upon the foundations of software reliability engineering, HIP-LLM defines LLM reliability as the probability of failure-free operation over a specified number of future tasks under a given Operational Profile (OP). HIP-LLM represents dependencies across (sub-)domains hierarchically, enabling multi-level inference from subdomain to system-level reliability. HIP-LLM embeds imprecise priors to capture epistemic uncertainty and incorporates OPs to reflect usage contexts. It derives posterior reliability envelopes that quantify uncertainty across priors and data. Experiments on multiple benchmark datasets demonstrate that HIP-LLM offers a more accurate and standardized reliability characterization than existing benchmark and state-of-the-art approaches. A publicly accessible repository of HIP-LLM is provided.
@@
Large Language Model; Reliability; Hierarchical Bayesian Inference; Operational Profile; Epistemic Uncertainty; Imprecise Probability
journal: Reliability Engineering & System Safety
@@ 1 Introduction
To bridge the five research gaps collectively, we propose HIP-LLM, a hierarchical imprecise probability approach to assess LLM reliability. HIP-LLM devises a more rigorous and versatile assessment with the following key features. First, similar to [31], it structures the model into independent high-level domains, with each domain further divided into dependent subdomains. In this way, reliability can be evaluated at multiple levels of detail while preserving statistical dependencies, so that information from one subdomain contributes to the inference about other subdomains within the same domain. We then incorporate OPs as operational weights at each level of the hierarchy (subdomain $\rightarrow$ domain $\rightarrow$ general-purpose LLM) to reflect the dynamic operational use of subdomains and domains in practice. In contrast to [31], which relies on non-informative priors, HIP-LLM enables the incorporation of informative prior knowledge. Since selecting a single prior in Bayesian inference is often controversial due to its encoding of epistemic uncertainty, we adopt Imprecise Probability [36, 37] that represents uncertainty about the prior itself without committing to a single distribution. Consequently, reliability metrics (defined in terms of both failure probability and the probability of failure-free runs in future operations) are expressed through posterior distribution envelopes at different levels of the hierarchy, reflecting both uncertainties from the data and the prior knowledge.
@@ 2.1 Software Reliability
The probability of failure-free software operation for a specified period of time in a specified environment
@@
There are some caveats regarding the term “inputs” in the definition of OP, particularly when considering factors such as software memory and dependencies [^4] In practice, because software systems may possess memory, their success/failure often depends on entire sequences of inputs rather than isolated ones. Accordingly, “inputs” may denote a complete stimulus, which may encompass the full sequence of interactions within a task, or even the cumulative sequence of inputs since the system was last reinitialized.. We refer readers to [23] for a detailed discussion.
In the context of LLMs, OPs serve an analogous role [27]: they specify the distribution of task types (e.g., factual question answering, coding, summarization) and their relative importance in deployment. Incorporating OPs into LLM reliability analysis ensures that evaluation metrics capture not just aggregate benchmark scores, but also how well a model performs across the realistic mix of tasks it is expected to face.
@@ 2.3 Statistical Inference for Evaluating LLMs
Despite these advances, important limitations remain. Neither Miller’s frequentist framework nor HiBayES explicitly defines/models the OP (even though both implicitly assume that the data represent it). Moreover, neither approach can effectively incorporate (potentially imprecise) prior knowledge: Miller’s work is restricted to classical frequentist inference, while HiBayES relies on non-informative priors. Both also focus on the narrower notion of failure probability as reliability, overlooking the more general and standardized reliability definition in terms of the probability of failure-free runs over future operations. The latter requires predictive inference that accounts for the propagation of uncertainties into the future: a more demanding but practical reliability claim [33, 34].
@@ 3.1 Problem Formulation
Let $\mathcal{X}$ denote the input-space of all possible tasks for a given LLM, and let $\pi$ be the OP, i.e., a probability distribution over $\mathcal{X}$ that reflects the likelihood of encountering each task $x\in\mathcal{X}$ in practice. Consider a sequence of $n\geq 1$ tasks independently and identically distributed (i.i.d.) according to the OP $\pi$, let $\mathbb{I}(x_{t})\in\{0,1\}$ indicate success (1) or failure (0) on the $t$-th task $x_{t}$, then the LLM reliability is:
@@
Intuitively, $R(n,\pi)$ represents the probability that the LLM will operate failure-free across the next $n$ i.i.d. tasks according to the OP $\pi$. Importantly, we make the following notes:
  1. 1.
    A key assumption of the above definition is the task failures/success are i.i.d. Bernoulli trials. Such modeling is not uncommon in reliability modeling, especially for critical on-demand systems. A typical justification is when demands are rare, and the states/memory of the software and its operational environment are effectively “reset” in-between [23, 71]. In the context of LLMs, we carefully define our reliability metric in terms of i.i.d “tasks”, rather than individual prompts which are often contextually dependent. A task may consist of a sequence of related prompts aimed at achieving a single task goal. We acknowledge that modern LLMs (e.g., ChatGPT) typically retain chat history as contextual memory, which can violate the independence assumption between tasks. However, most LLMs also provide the option to start a new chat session (for a new single task) without any memory and history from previous chat sessions, aligning with the assumption of “resetting” the LLM.
  2. 2.
    While $R(n,\pi)$ is the general form of future reliability of processing $n$ tasks, the special case $1-R(1,\pi)$ (where $n=1$) represents the failure probability [33, 34, 35] that studied by, e.g., [31].
  3. 3.
@@
To rigorously assess the formally defined LLM reliability, coping with the aforementioned notes, the next subsection introduces our proposed solution HIP-LLM. It models the LLM as a hierarchical structure consisting of independent domains, each containing statistically dependent subdomains (cf. Fig. 1). Here the modeled dependencies and independencies represent the epistemic structure of our hierarchical Bayesian model. That is, observing failures in one (sub-)domain may or may not update our beliefs about the reliability of other (sub-)domains. One possible example of justification [^8] Cf. Section 5 for discussions on the validity of this hierarchal dependency assumption. is: We model coding and law as independent because they rely on distinct competencies of the LLM—coding on formal, symbolic reasoning and syntax manipulation, and law on narrative understanding and normative interpretation. Since these skills draw from largely separate representations and training data, failures in one domain provide little information about failures in the other. In Bayesian terms, their failure probabilities can be treated as a priori independent parameters, reflecting separate latent skill dimensions of the model. On the other hand, sub-domains failure probabilities are modeled as dependent parameters given their shared LLM competencies.
@@
Our goal is to infer the posterior distributions of future reliability (and its special case the failure probability), at the subdomain, domain, and general-purpose LLM levels, based on observed correct responses from tasks within each subdomain.
@@
Figure 1: Schematic representation of the hierarchical LLM, domain, and subdomain structure for reliability estimation fo M LLM models.
@@ 3.2 Proposed Solution
Fig. 2 presents a detailed view of the hierarchical structure (subdomains $\rightarrow$ domains $\rightarrow$ LLM), with assumed priors and parameters. We assume that subdomain reliabilities within a domain are dependent through a shared prior, and that domain reliabilities are aggregated from their subdomains according to task-specific operational profiles. Our goal is to construct a principled hierarchical Bayesian model that supports information sharing across dependent subdomains through partial pooling (Sec. 3.2.1) and uncertainty quantification via Imprecise Probability (Sec. 3.2.2).
@@
Figure 2: Hierarchical structure with independent domains and dependent subdomains. For readability, parameters and priors are shown only for one subdomain $S_{ij}$ and its parent domain $D_{i}$ under a representative $LLM_{k}$; the remaining subdomains, domains ($i=1,\dots,m$), and LLMs $(k=1,\dots,M)$ are identical, differing only in their indices.
@@ 3.2.3 Theorems
However, for later Theorems closed-form densities do not exist because $p_{i}=\sum_{j}\Omega_{ij}\theta_{ij}$ and $p_{L}=\sum_{i}W_{i}p_{i}$ are weighted sums of dependent random variables—the distribution of a sum of Beta random variables has no analytical form except in trivial cases [^15] Assuming the weights are constants. When the weights are modeled as random variables with their own probability distributions, the same problem persists (if not harder).. Computing such densities would require intractable multi-dimensional integrals. The Cumulative Distribution Function (CDF) formulation sidesteps this problem: $F_{p_{i}}(t\mid C_{i},h_{i})=\int_{0}^{1}\int_{0}^{\infty}F_{p_{i}}(t\mid\mu_{i},\nu_{i},C_{i})\,Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i}$ only requires a two-dimensional integral over $(\mu_{i},\nu_{i})$, where the conditional CDF can be computed via Monte Carlo sampling of independent Betas. Since CDFs provide all necessary information for practical reliability assessment (probabilities, quantiles, expectations), they are the natural representation when densities are unavailable.
@@
Then, for any $h_{i}\in\mathcal{A}_{i}$, the posterior distribution of $p_{i}$ is characterized by its cumulative distribution function:
@@
Then, for any $\mathcal{H}\in\mathcal{A}_{\text{LLM}}$, the posterior distribution of $p_{L}$ is characterized by its cumulative distribution function:
@@
For any $h_{i}\in\mathcal{A}_{i}$, the posterior cumulative distribution function of $R_{ij}(n_{F})$ is:
$F_{R_{ij}(n_{F})}(r\mid C_{i},h_{i})=\Pr(\theta_{ij}^{n_{F}}\leq r\mid C_{i},h_{i})=\int_{0}^{r^{1/n_{F}}}Pr(\theta_{ij}\mid C_{i},h_{i})\,d\theta_{ij},$
@@
$\underline{F}_{R_{ij}(n_{F})}(r\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}F_{R_{ij}(n_{F})}(r\mid C_{i},h_{i}),$
$\overline{F}_{R_{ij}(n_{F})}(r\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}F_{R_{ij}(n_{F})}(r\mid C_{i},h_{i})$
@@
$F_{R_{i}(n_{F})}(r\mid C_{i},h_{i})=\Pr\left(\left[\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}\right]^{n_{F}}\leq r\,\bigg|\,C_{i},h_{i}\right),$
@@
$\underline{F}_{R_{i}(n_{F})}(r\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}F_{R_{i}(n_{F})}(r\mid C_{i},h_{i}),$
$\overline{F}_{R_{i}(n_{F})}(r\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}F_{R_{i}(n_{F})}(r\mid C_{i},h_{i})$
@@
$F_{R_{L}(n_{F})}(r\mid\text{data},\mathcal{H})=\Pr\left(\left[\sum_{i=1}^{m}W_{i}p_{i}\right]^{n_{F}}\leq r\,\bigg|\,\text{data},\mathcal{H}\right),$
@@
$\underline{F}_{R_{L}(n_{F})}(r\mid\text{data})=\inf_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{R_{L}(n_{F})}(r\mid\text{data},\mathcal{H}),$
$\overline{F}_{R_{L}(n_{F})}(r\mid\text{data})=\sup_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{R_{L}(n_{F})}(r\mid\text{data},\mathcal{H})$

### v2

# A Hierarchical Imprecise Probability Approach to Reliability Assessment of Large Language Models

@@ Abstract
Large Language Models (LLMs) are increasingly deployed across diverse domains, raising the need for rigorous reliability assessment methods. Existing benchmark-based evaluations primarily offer descriptive statistics of model accuracy over datasets, providing limited insight into the probabilistic behavior of LLMs under real operational conditions. This paper introduces HIP-LLM, a Hierarchical Imprecise Probability framework for modeling and inferring LLM reliability. Building upon the foundations of software reliability engineering, HIP-LLM defines LLM reliability as the probability of failure-free operation over a specified number of future tasks under a given Operational Profile (OP). HIP-LLM represents dependencies across (sub-)domains hierarchically, enabling multi-level inference from subdomain to system-level reliability. HIP-LLM embeds imprecise priors to capture epistemic uncertainty and incorporates OPs to reflect usage contexts. It derives posterior reliability envelopes that quantify uncertainty across priors and data. Experiments on multiple benchmark datasets demonstrate that HIP-LLM offers a more nuanced and standardized reliability characterization than existing benchmark and state-of-the-art approaches. A publicly accessible repository of HIP-LLM is provided.
@@
Large Language Model; Software Reliability; Hierarchical Bayesian Inference; Operational Profile; Epistemic Uncertainty; Imprecise Probability
@@ 1 Introduction
To bridge the five research gaps collectively, we propose HIP-LLM, a hierarchical imprecise probability approach to assess LLM reliability. HIP-LLM devises a more rigorous and versatile assessment with the following key features. First, similar to [43], it structures the model into independent high-level domains, with each domain further divided into dependent subdomains. In this way, reliability can be evaluated at multiple levels of detail while preserving statistical dependencies, so that information from one subdomain contributes to the inference about other subdomains within the same domain. We then incorporate OPs as operational weights at each level of the hierarchy (subdomain $\rightarrow$ domain $\rightarrow$ general-purpose LLM) to reflect the dynamic operational use of subdomains and domains in practice. We assume tasks constitute independent and identically distributed (i.i.d.) trials drawn from the specified OP. This assumption is appropriate for reset or single-task scenarios, in which each task is executed in a fresh session without shared context or memory—settings commonly encountered in offline evaluation like benchmarking [^3] In contrast, long-context or agentic workflows, where task outcomes are sequentially dependent through memory or tool use, violate the i.i.d. assumption and require fundamentally different modeling approaches to our HIP-LLM. We provide more discussions later in Remark 1 and Section 5. Accordingly, HIP-LLM is intended for LLM usage scenarios in which each task is executed in a new chat session without retaining memory from previous sessions, thereby aligning with the assumption of a “reset” LLM.
In contrast to [43], which relies on non-informative priors, HIP-LLM enables the incorporation of informative prior knowledge. Since selecting a single prior in Bayesian inference is often controversial due to its encoding of epistemic uncertainty, we adopt Imprecise Probability [4, 68] that represents uncertainty about the prior itself without committing to a single distribution. Consequently, reliability metrics (defined in terms of both failure probability and the probability of failure-free runs in future operations) are expressed through posterior distribution envelopes at different levels of the hierarchy, reflecting both uncertainties from the data and the prior knowledge.
@@ 2.1 Software Reliability
The probability of failure-free software operation for a specified period of time in a specified environment.
@@
There are some caveats regarding the term “inputs” in the definition of OP, particularly when considering factors such as software memory and dependencies [^5] In practice, because software systems may possess memory, their success/failure often depends on entire sequences of inputs rather than isolated ones. Accordingly, “inputs” may denote a complete stimulus, which may encompass the full sequence of interactions within a task, or even the cumulative sequence of inputs since the system was last reinitialized.. We refer readers to [63] for a detailed discussion. For LLMs, OPs serve an analogous role [15]: they specify the distribution of task types (e.g., factual question answering, coding, summarization) and their relative importance in deployment. Incorporating OPs into LLM reliability analysis ensures that evaluation metrics capture not just aggregate benchmark scores, but also how well a model performs across the realistic mix of tasks it is expected to face.
@@ 2.3 Statistical Inference for Evaluating LLMs
Despite these advances, important limitations remain. Neither Miller’s frequentist framework nor HiBayES explicitly defines/models the OP (even though both implicitly assume that the data represent it). Moreover, neither approach can effectively incorporate (potentially imprecise) prior knowledge: Miller’s work is restricted to classical frequentist inference, while HiBayES relies on non-informative priors. Both also focus on the narrower notion of failure probability as reliability, overlooking the more general and standardized reliability definition in terms of the probability of failure-free runs over future operations. The latter requires predictive inference that accounts for the propagation of uncertainties into the future: a more demanding but practical reliability claim [64, 8]. Miller accounts for clustered dependence through variance corrections, but does not introduce hierarchical latent variables or multi-level generative models as in HiBayES or HIP-LLM. Table 1 summarizes the key features and provides a comparative overview of the three methods (HIP-LLM, HiBayES [43], and Miller [46]).
Table 1: Comparison of LLM evaluation frameworks highlighting differences in reliability definition, operational profile modeling, and uncertainty treatment
Aspect Miller [46] HiBayES [43] HIP-LLM
Statistical framework Frequentist Bayesian Bayesian
Primary goal Uncertainty-aware benchmark reporting Uncertainty-aware, hierarchical accuracy estimation Uncertainty-aware reliability assessment under OP with embedded priors
Metric under estimation Failure probability (per task) Failure probability (per task) Reliability (probability of failure-free future tasks)
OP Implicit (dataset assumed representative) Implicit (dataset assumed representative) Explicit, multi-level OPs
Hierarchical structure None LLM, domains, subdomains LLM, domains, subdomains
Dependence handling Clustered standard errors Partial pooling Partial pooling
Prior specification None Uniform precise priors Informative imprecise priors (credal sets)
Output Confidence intervals Single posterior distribution Posterior envelopes
@@ 3.1 Problem Formulation
Let $\mathcal{X}$ denote the input-space of all possible tasks for a given LLM, and let $\pi$ be the OP, i.e., a probability distribution over $\mathcal{X}$ that reflects the likelihood of encountering each task $x\in\mathcal{X}$ in practice. Consider a sequence of $n\geq 1$ tasks i.i.d. according to the OP $\pi$, let $\mathbb{I}(x_{t})\in\{0,1\}$ indicate success (1) or failure (0) on the $t$-th task $x_{t}$, then the LLM reliability is:
@@
Intuitively, $R(n,\pi)$ represents the probability that the LLM will operate failure-free across the next $n$ i.i.d. tasks according to the OP $\pi$. Importantly, we make the following remarks:
Remark 1 (The i.i.d. assumption vs. contextual memory).
A key assumption of the above definition is the task failures/success are i.i.d. Bernoulli trials. Such modeling is not uncommon in reliability modeling, especially for critical on-demand systems. A typical justification is when demands are rare, and the states/memory of the software and its operational environment are effectively “reset” in-between [63, 58]. In the context of LLMs, we carefully define our reliability metric in terms of i.i.d “tasks”, rather than individual prompts which are often contextually dependent. A task may consist of a sequence of related prompts aimed at achieving a single task goal. We acknowledge that modern LLMs (e.g., ChatGPT) typically retain chat history as contextual memory, which can violate the i.i.d. assumption between tasks. However, most LLMs provide the option to start a new chat session (for a new single task) without any memory and history from previous chat sessions, aligning with the assumption of “resetting” the LLM. We note such “reset” settings are commonly encountered in LLM offline evaluation like benchmarking [13]. In contrast, long-context or agentic workflows, where tasks are sequentially dependent through memory or tool use, violate this assumption and require alternative stateful reliability models which is out of the scope of this paper.
Remark 2 (Failure probability vs. future reliability).
While the reliability metric $R(n,\pi)$ is the general form of future reliability of processing $n$ tasks, the special case $1-R(1,\pi)$ (where $n=1$) represents the failure probability [64, 8, 85] that studied by, e.g., [46, 43]. Accordingly, throughout this paper, we use the term failure probability to denote $1-R(1,\pi)$ and failure-free runs (of $n$ future tasks) to denote $R(n,\pi)$, in order to distinguish these two notions. The more general term reliability is used to refer to either quantity when the intended meaning is clear from the surrounding context.
Remark 3 (General purpose reliability vs. domain-specific reliability).
@@
Remark 4 (Binary failure vs. non-binary scoring).
In traditional software reliability engineering, failures are naturally defined as binary events, success or failure, according to an explicit system specification. Similarly, most LLM benchmarks introduce task-specific specifications, implemented via automated evaluators or human annotations, which yield binary outcomes for scoring each prompted task. Consistent with this established practice, we model LLM reliability based on binary failure in Def. 4. However, unlike traditional software, what constitutes a “failure” for an LLM can be inherently subjective and domain dependent, and external scoring mechanisms may themselves be noisy or inconsistent. Accordingly, reliability in HIP-LLM in Def. 4 is formalized conditional on a specified failure definition and scoring process, rather than as an absolute, model-intrinsic property. Explicitly modeling scoring uncertainty or alternative failure definitions remains an important direction for future work.
To do statistical inference for the reliability metric defined in Eq. (1) with assumptions in aforementioned remarks, a simplified “textbook” Bayesian model would be the Beta-Binominal one (which also used as one of the baselines in our experiments). This Beta-Binomial estimator applies to a single domain with precise prior knowledge. Let $\theta:=\Pr_{x\sim\pi}(I(x)=1)$ denote the (unknown) probability of success on a random task drawn from the OP $\pi$. Given $N$ i.i.d. evaluated tasks with $C$ successes and NCN-C failures, assume a Binomial likelihood
$C,N\mid\theta\sim\mathrm{Binomial}(N,\theta)=\theta^{C}(1-\theta)^{N-C},$
and a prior $\Pr(\theta)$. By Bayes’ rule, the posterior distribution of $\theta$ is
$\Pr(\theta\mid C,N)=\frac{\theta^{C}(1-\theta)^{N-C}\Pr(\theta)}{\int_{0}^{1}\theta^{C}(1-\theta)^{N-C}\Pr(\theta)\,d\theta}.$ (2)
Similarly for the future reliability of passing $n_{F}$ tasks:
$\Pr(R(n_{F},\pi)\mid C,N)=\frac{\theta^{n_{F}}\theta^{C}(1-\theta)^{N-C}\Pr(\theta)}{\int_{0}^{1}\theta^{C}(1-\theta)^{N-C}\Pr(\theta)}$ (3)
If $\Pr(\theta)=\mathrm{Beta}(\alpha,\beta)$, conjugacy yields the closed-form posterior
$\theta\mid C,N\sim\mathrm{Beta}(\alpha+C,\;\beta+N-C).$
For future reliability, thanks to the conjugacy again, the posterior mean reliability for $n_{F}$ future tasks is therefore:
$\mathbb{E}[R(n_{F},\pi)\mid\mathcal{D}]=\mathbb{E}[\theta^{n_{F}}\mid\mathcal{D}]=\frac{B(\alpha+C+n_{F},\;\beta+N-C)}{B(\alpha+C,\;\beta+N-C)},$
where $B(\cdot,\cdot)$ denotes the Beta function. Similarly, the posterior PDF and CDF can also be derived and we omit them for brevity.
To more rigorously assess the formally defined LLM reliability, coping with the aforementioned Remarks and Gaps, the next subsection introduces our proposed solution HIP-LLM. It models the LLM as a hierarchical structure consisting of independent domains, each containing statistically dependent subdomains (cf. Fig. 1).
Remark 5 (The need of hierarchical modeling on (sub-)domain (in-)dependencies).
@@
The modeled dependencies and independencies represent the epistemic structure of our hierarchical Bayesian model. That is, observing failures in one (sub-)domain may or may not update our beliefs about the reliability of other (sub-)domains. One possible example of justification [^9] Cf. Section 5 for discussions on the validity of this hierarchal dependency assumption. is: We model coding and law as independent because they rely on distinct competencies of the LLM—coding on formal, symbolic reasoning and syntax manipulation, and law on narrative understanding and normative interpretation. Since these skills draw from largely separate representations and training data, failures in one domain provide little information about failures in the other. In Bayesian terms, their failure probabilities can be treated as a priori independent parameters, reflecting separate latent skill dimensions of the model. On the other hand, sub-domains failure probabilities are modeled as dependent parameters given their shared LLM competencies.
@@
Our goal is to infer the posterior distributions of future reliability (and its special case, the failure probability), at the subdomain, domain, and general-purpose LLM levels, based on observed correct responses from tasks within each subdomain.
@@
Figure 1: Schematic representation of the hierarchical LLM, domain, and subdomain structure for reliability estimation for M LLM models. Rectangles indicate independent components, while ovals indicate dependent components within the hierarchy.
@@ 3.2 Proposed Solution
Fig. 2 presents a detailed view of the hierarchical structure (subdomains $\rightarrow$ domains $\rightarrow$ LLM), with assumed priors and parameters. We assume that subdomain reliabilities within a domain are dependent through a shared prior, and that domain reliabilities are aggregated from their subdomains according to task-specific OPs. Our goal is to construct a principled hierarchical Bayesian model that supports information sharing across dependent subdomains through partial pooling (Sec. 3.2.1) and uncertainty quantification via Imprecise Probability (Sec. 3.2.2).
@@
Figure 2: Hierarchical structure with independent domains and dependent subdomains. For readability, parameters and priors are shown only for one subdomain $S_{ij}$ and its parent domain $D_{i}$ under a representative $LLM_{k}$; the remaining subdomains, domains ($i=1,\dots,m$), and LLMs $(k=1,\dots,M)$ are identical, differing only in their indices. Rectangles indicate independent components, while ovals indicate dependent components within the hierarchy.
@@ 3.2.2 Uncertainty Handling via Imprecise Probability
Here and in the following, $t\in[0,1]$ denotes a generic probability threshold at which the CDF of the corresponding (non-failure probability or reliability) random variable is evaluated.
@@ 3.2.3 Theorems
However, for later Theorems closed-form densities do not exist because $p_{i}=\sum_{j}\Omega_{ij}\theta_{ij}$ and $p_{L}=\sum_{i}W_{i}p_{i}$ are weighted sums of dependent random variables—the distribution of a sum of Beta random variables has no analytical form except in trivial cases [^16] Assuming the weights are constants. When the weights are modeled as random variables with their own probability distributions, the same problem persists (if not harder).. Computing such densities would require intractable multi-dimensional integrals. The CDF formulation sidesteps this problem: $F_{p_{i}}(t\mid C_{i},h_{i})=\int_{0}^{1}\int_{0}^{\infty}F_{p_{i}}(t\mid\mu_{i},\nu_{i},C_{i})\,Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i}$ only requires a two-dimensional integral over $(\mu_{i},\nu_{i})$, where the conditional CDF can be computed via Monte Carlo sampling of independent Betas. Since CDFs provide all necessary information for practical reliability assessment (probabilities, quantiles, expectations), they are the natural representation when densities are unavailable.
@@
Then, for any $h_{i}\in\mathcal{A}_{i}$, the posterior distribution of $p_{i}$ is characterized by its CDF:
@@
Then, for any $\mathcal{H}\in\mathcal{A}_{\text{LLM}}$, the posterior distribution of $p_{L}$ is characterized by its CDF:
@@
For any $h_{i}\in\mathcal{A}_{i}$, the posterior CDF of $R_{ij}(n_{F})$ is:
$F_{R_{ij}(n_{F})}(t\mid C_{i},h_{i})=\Pr(\theta_{ij}^{n_{F}}\leq t\mid C_{i},h_{i})=\int_{0}^{r^{1/n_{F}}}Pr(\theta_{ij}\mid C_{i},h_{i})\,d\theta_{ij},$
@@
$\underline{F}_{R_{ij}(n_{F})}(t\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}F_{R_{ij}(n_{F})}(t\mid C_{i},h_{i}),$
$\overline{F}_{R_{ij}(n_{F})}(t\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}F_{R_{ij}(n_{F})}(t\mid C_{i},h_{i})$
@@
$F_{R_{i}(n_{F})}(t\mid C_{i},h_{i})=\Pr\left(\left[\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}\right]^{n_{F}}\leq t\,\bigg|\,C_{i},h_{i}\right),$
@@
$\underline{F}_{R_{i}(n_{F})}(t\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}F_{R_{i}(n_{F})}(t\mid C_{i},h_{i}),$
$\overline{F}_{R_{i}(n_{F})}(t\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}F_{R_{i}(n_{F})}(t\mid C_{i},h_{i})$
@@
$F_{R_{L}(n_{F})}(t\mid\text{data},\mathcal{H})=\Pr\left(\left[\sum_{i=1}^{m}W_{i}p_{i}\right]^{n_{F}}\leq t\,\bigg|\,\text{data},\mathcal{H}\right),$
@@
$\underline{F}_{R_{L}(n_{F})}(t\mid\text{data})=\inf_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{R_{L}(n_{F})}(t\mid\text{data},\mathcal{H}),$
$\overline{F}_{R_{L}(n_{F})}(t\mid\text{data})=\sup_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{R_{L}(n_{F})}(t\mid\text{data},\mathcal{H})$
@@
Remark 6 (Translate vague prior knowledge to hyperparameters).
In HIP-LLM, hyperparameters are not tuning constants but variables used to represent uncertainty about prior knowledge. At the lowest level, $\theta_{ij}$ indicates “how reliable” (i.e., the probability that the model succeeds on a randomly drawn task from the subdomain) of $S_{ij}$. Subdomains within the same domain $D_{i}$ are assumed to be related, and this dependence is captured by assuming that their non-failure probabilities are drawn from a shared Beta distribution governed by two domain-level parameters: $\mu_{i}$ and $\nu_{i}$. Here, $\mu_{i}$ represents the expected reliability of the LLM in domain $D_{i}$, while $\nu_{i}$ represents how confident we are in this expectation, or equivalently, how strongly subdomain reliabilities are expected to cluster around $\mu_{i}$. Since assessors typically do not know the exact values of $\mu_{i}$ and $\nu_{i}$, HIP-LLM places distributions on them: $\mu_{i}\sim\mathrm{Beta}(a_{i},b_{i})$ and $\nu_{i}\sim\mathrm{Gamma}(c_{i},d_{i})$. The hyperparameters $(a_{i},b_{i},c_{i},d_{i})$ therefore encode high-level and (likely) vague/imperfect expert beliefs rather than performance data. For example, suppose an assessor believes that the LLM’s reasoning-domain reliability is likely between $0.8$ and $0.9$, but with only moderate confidence corresponding to roughly 3030100100 observations (either from previous-use/historical-data or structured expert elicitation experiments based on hypothetical usage scenarios). This belief can be expressed by choosing $a_{i}\in[8,18]$ and $b_{i}\in[2,4]$, which yields prior means $\mathbb{E}[\mu_{i}]\in[0.8,0.9]$, and by choosing [^18] The numerical ranges are obtained using the standard interpretations of the Beta and Gamma distributions. For the domain-level mean reliability $\mu_{i}\sim\mathrm{Beta}(a_{i},b_{i})$, the prior mean is $\mathbb{E}[\mu_{i}]=a_{i}/(a_{i}+b_{i})$. Choosing $a_{i}$ and $b_{i}$ such that this ratio lies between $0.8$ and $0.9$ ensures that the assessor’s belief about the expected reliability is respected (e.g., $8/(8+2)=0.8$ and $18/(18+2)=0.9$). The exact values of $(a_{i},b_{i})$ are not unique; the interval reflects uncertainty about the precise prior shape. For the confidence parameter $\nu_{i}\sim\mathrm{Gamma}(c_{i},d_{i})$ (shape–rate form), the prior mean is $\mathbb{E}[\nu_{i}]=c_{i}/d_{i}$, which can be interpreted as an equivalent number of pseudo-observations supporting the prior belief. Setting $d_{i}=0.1$ and letting $c_{i}$ vary between $3$ and $10$ yields $\mathbb{E}[\nu_{i}]\in[30,100]$, matching the assessor’s statement of moderate confidence. $c_{i}\in[3,10]$ and $d_{i}=0.1$, which gives $\mathbb{E}[\nu_{i}]\in[30,100]$. By allowing these hyperparameters to vary within intervals rather than fixing them to single values, HIP-LLM captures realistic epistemic uncertainty (following the imprecise probability framework [4, 68] designed to facilitate elicitation of imperfect/vague prior knowledge) and propagates this uncertainty through the hierarchy, yielding posterior reliability bounds rather than overconfident point estimates.